OpenStudy (anonymous):

"If two lines intersect, then their intersection is A. a line." B. at least one point." C. one point." D. a midpoint."

OpenStudy (anonymous):

Is C the correct answer?

ganeshie8 (ganeshie8):

C is correct

ganeshie8 (ganeshie8):



ganeshie8 (ganeshie8):

that case we call overlapping lines

ganeshie8 (ganeshie8):

we dont call two overlapping lines as intersecting

ganeshie8 (ganeshie8):

intersecting and overlapping are two different terms in geometry

ganeshie8 (ganeshie8):

when lines overlap, you get infinite solutions. when lines intersect, you get oly one solution

ganeshie8 (ganeshie8):

ha :) i get you... in highschool level we just teach students with these definitions... cuz we need to maintain some clarity about terminology... they will see things later clearly :)

ganeshie8 (ganeshie8):

i dont think so, see how many questions RH and myself worked together in the past

ganeshie8 (ganeshie8):

interesting how can anyone prove 1+1=2 ha ?

ganeshie8 (ganeshie8):

I bet u will :)

ganeshie8 (ganeshie8):

in general 9 is never 3 and 1+1 is never 1

ganeshie8 (ganeshie8):

ha

ganeshie8 (ganeshie8):

1+1 = 1 in bloolean algebra, cuz + is OR operator is it

ganeshie8 (ganeshie8):

1+1 = 10 right ?

ganeshie8 (ganeshie8):

I see what you mean, the more we learn the more paranoid we get ;p

ganeshie8 (ganeshie8):

did you read the book "Only the paranoid survives"

ganeshie8 (ganeshie8):

ha ok :) cya have a great day

OpenStudy (loser66):

thanks , cya, but I have to delete the conversation, it shows my dummy brain, said something nonsense